OpenStudy (anonymous):

Solve. 4x^2 + 8x = 252 x = 7 and x = 9 x = -7 and x = 9 x = -7 and x = -9 x = 7 and x = -9

myininaya (myininaya):

i think it is easy to divide everything by 4 first \[x^2+2x=63 => x^2+2x-63=0\]

myininaya (myininaya):

now we can easily factor this

myininaya (myininaya):

\[(x+9)(x-7)=0\] since 9(-7) is -63 and 9+(-7)=2

myininaya (myininaya):

now set both factors =0 and solve for x